# arXiv:2603.02442v1[math.FA] 02 Mar 2026

### 02 Mar 2026 v1

@@ 4. The $H^{\infty}$ space
Let $C_{w,\varphi}$ be a weighted composition operator on $H^{\infty}$. Then $C_{w,\varphi}$ is mean Li-Yorke chaotic if there is an increasing sequence $(N_{k})_{k\in\mathbb{N}}$ of positive integers such that $\lim_{k\to\infty}\dfrac{1}{N_{k}}\sum_{j=1}^{N_{k}}\|w^{(j)}\|_{\infty}=0$, and $\sup_{N\in\mathbb{N}}\dfrac{1}{N}\sum_{j=1}^{N}\|w^{(j)}\|_{\infty}=\infty.$
5. Example
We now present an example illustrating the results established in Section 3 for weighted composition operators on Hardy spaces.
@@ 5. Example
It was shown in [8] that the point spectrum of the composition operator $C_{\varphi_{a}}$ on the Hardy space $H^{2}$ contains the annulus $\{\lambda\in\mathbb{C}:0<|\lambda|<a^{-1/2}\}$. This allows us to select $k,\lambda\in\mathbb{C}$ and $f\in H^{2}$ satisfying the following conditions simultaneously:
  • $1<|k|<a^{-1/2}$,
  • $a^{1/2}<|\lambda|<1$,
  • $|k\cdot\lambda|>1$,
  • $C_{\varphi_{a}}f=kf$.
Since $C_{\varphi_{a^{n}}}w_{\lambda}-\lambda=\lambda a^{n}(z-1)$, the sequence $C_{\varphi_{a^{n}}}w_{\lambda}$ converges to the constant function $\lambda$ in the $H^{\infty}$-norm. Because $|k\cdot\lambda|>1$, we can find $n_{0}\in\mathbb{N}$ and $\gamma<|\lambda|$ with $|\gamma\cdot k|>1$, such that $|C_{\varphi_{a^{n}}}w_{\lambda}(z)|>\gamma$ for all $z\in\mathbb{D}$ and all $n>n_{0}$.
@@
$|(C_{w_{\lambda},\varphi_{a}})^{n}f(z)|\geq|k\cdot\gamma|^{n-n_{0}}\,|(C_{w_{\lambda},\varphi_{a}})^{n_{0}}f(z)|,$
@@
$\|(C_{w_{\lambda},\varphi_{a}})^{n}f\|_{2}\geq|k\cdot\gamma|^{n-n_{0}}\,\|(C_{w_{\lambda},\varphi_{a}})^{n_{0}}f\|_{2}.$
As $|k\cdot\gamma|>1$, the right-hand side grows without bound, showing that $C_{w_{\lambda},\varphi_{a}}$ is not power bounded. Theorem 4 then implies that $C_{w_{\lambda},\varphi_{a}}$ is Li–Yorke chaotic on $H^{2}$. Finally, since $H^{2}\subseteq H^{p}$ for $1\leq p\leq 2$, the same argument carries over to show that $C_{w_{\lambda},\varphi_{a}}$ is Li–Yorke chaotic on $H^{p}$ for all $1\leq p\leq 2$.

### 13 Mar 2026 v2

@@ 4. The $H^{\infty}$ space
Let $C_{w,\varphi}$ be a weighted composition operator on $H^{\infty}$. Then $C_{w,\varphi}$ is mean Li-Yorke chaotic if there is an increasing sequence $(N_{k})_{k\in\mathbb{N}}$ of positive integers such that
$\lim_{k\to\infty}\dfrac{1}{N_{k}}\sum_{j=1}^{N_{k}}\|w^{(j)}\|_{\infty}=0\quad\text{and}\quad\displaystyle\sup_{N\in\mathbb{N}}\dfrac{1}{N}\sum_{j=1}^{N}\|w^{(j)}\|_{\infty}=\infty.$
5. Li-Yorke chaotic weighted compostion operators on $H^{p}$ and $A^{2}_{\beta}$
This section is devoted to examples. We present an example illustrating the results established in Section 3 for weighted composition operators on Hardy and Bergman spaces. On the other hand, we provide an example where the condition in Theorem 4 is not strictly necessary to obtain Li-Yorke chaos.
@@ 5. Li-Yorke chaotic weighted compostion operators on $H^{p}$ and $A^{2}_{\beta}$
Proposition 3.
Let $0<a<1$ and suppose that $a^{1/2}<|\lambda|<1$. Then the weighted composition operator $C_{w_{\lambda},\varphi_{a}}$ is densely Li–Yorke chaotic on $A^{2}_{\beta}$ and $H^{p}$ for $1\leq p\leq 2$.
Proof.
First, choose $\mu\in\mathbb{R}$ such that $1<\mu<a^{-1/2}$ and $|\mu\cdot\lambda|>1$. It was shown in [10] that the point spectrum of the composition operator $C_{\varphi_{a}}$ on the Hardy space $H^{2}$ contains the annulus $\{\mu\in\mathbb{C}:0<|\mu|<a^{-1/2}\}$. Then there exists $f_{\mu}\in H^{2}$ satisfying $C_{\varphi_{a}}f_{\mu}=\mu f_{\mu}$. Since $C_{\varphi_{a^{n}}}w_{\lambda}-\lambda=\lambda a^{n}(z-1)$, the sequence $C_{\varphi_{a^{n}}}w_{\lambda}$ converges to the constant function $\lambda$ in the $H^{\infty}$-norm. Because $|\mu\cdot\lambda|>1$, we can find $n_{0}\in\mathbb{N}$ and $\gamma<|\lambda|$ with $|\mu\cdot\gamma|>1$, such that $|C_{\varphi_{a^{n}}}w_{\lambda}(z)|>\gamma$ for all $z\in\mathbb{D}$ and all $n>n_{0}$.
@@
$|(C_{w_{\lambda},\varphi_{a}})^{n}f_{\mu}(z)|\geq|\mu\cdot\gamma|^{n-n_{0}}\,|(C_{w_{\lambda},\varphi_{a}})^{n_{0}}f_{\mu}(z)|,$
@@
$\|(C_{w_{\lambda},\varphi_{a}})^{n}f_{\mu}\|_{2}\geq|\mu\cdot\gamma|^{n-n_{0}}\,\|(C_{w_{\lambda},\varphi_{a}})^{n_{0}}f_{\mu}\|_{2}.$
As $|\mu\cdot\gamma|>1$, the right-hand side grows without bound, showing that $C_{w_{\lambda},\varphi_{a}}$ is not power bounded. Theorem 4 then implies that $C_{w_{\lambda},\varphi_{a}}$ is densely Li–Yorke chaotic on $H^{2}$. Finally, since $H^{2}\subseteq A^{2}_{\beta}$ and $H^{2}\subseteq H^{p}$ for $1\leq p\leq 2$, the same argument carries over to show that $C_{w_{\lambda},\varphi_{a}}$ is densely Li–Yorke chaotic on $A^{2}_{\beta}$ and $H^{p}$ for all $1\leq p\leq 2$.
In what follows, we show that the condition $\|w^{(n)}\|_{X}\to 0$ in Theorem 4 is not necessary for a weighted composition operator to be Li–Yorke chaotic.
Proposition 4.
Let $0<a<1$. Then the composition operator $C_{\varphi_{a}}$ is Li–Yorke chaotic on $H^{2}$ and $A^{2}_{\beta}$.
Proof.
In [10] it was shown that the functions $f_{s}(z)=(z-1)^{s}$ are eigenvectors of $C_{\varphi_{a}}$ and satisfy
$C_{\varphi_{a}}f_{s}=a^{s}f_{s},$
where $f_{s}\in H^{2}$ if and only if $\operatorname{Re}(s)>-\frac{1}{2}$. It is straightforward to verify that
  • $\|C_{\varphi_{a}}^{n}f_{s}\|_{2}\to 0$ as $n\to\infty$ if $\mathrm{Re}(s)>0$.
  • $\|C_{\varphi_{a}}^{n}f_{s}\|_{2}\to\infty$ as $n\to\infty$ if $\mathrm{Re}(s)<0$.
Let us check that, for each $k\in\mathbb{N}$, $f_{1/4}z^{k}\in\text{NS}(C_{\varphi_{a}})$. Indeed:
$\|C_{\varphi_{a}}^{n}f_{1/4}z^{k}\|_{2}$ $=\|C_{\varphi_{a}}^{n}f_{1/4}C_{\varphi_{a}}^{n}z^{k}\|_{2}=\|a^{n/4}f_{1/4}C_{\varphi_{a}}^{n}z^{k}\|_{2}$
$\leq a^{n/4}\|f_{1/4}\|_{\infty}\|C_{\varphi_{a^{n}}}z^{k}\|_{\infty}$
$=a^{n/4}\|f_{1/4}\|_{\infty}\|(a^{n}z+1-a^{n})^{k}\|_{\infty}$
$\leq a^{n/4}\|f_{1/4}\|_{\infty}(a^{n}+1)^{k}.$
Therefore, $\|C_{\varphi_{a}}^{n}f_{1/4}z^{k}\|_{2}\rightarrow 0$ as $n\rightarrow\infty$. Since $f_{1/4}\in H^{\infty}$, it follows that
$f_{1/4}H^{2}\subseteq\mathscr{X}=\overline{\operatorname{span}}(\mathrm{NS}(C_{\varphi_{a}})).$
In particular, $f_{-1/12}=f_{1/4}f_{-1/3}\in\mathscr{X}$, and since
$\left\|C_{\varphi_{a}}^{\,n}f_{-1/12}\right\|_{2}\to\infty\quad\text{as }n\to\infty,$
we conclude that $C_{\varphi_{a}}$ is Li–Yorke chaotic by Proposition 2. The same argument holds for $A^{2}_{\beta}$. ∎
Recall that $\mathbb{C}_{+}:=\{z\in\mathbb{C}:\text{Re}(z)>0\}$ represents the open right half-plane. In [1, Theorem 10] it was proved that no composition operator with linear fractional symbol on the Hardy space $H^{2}(\mathbb{C}_{+})$ can be Li-Yorke chaotic. We now exhibit a family of Li–Yorke chaotic operators which are a constant multiple of composition operators on $H^{2}(\mathbb{C}_{+})$.
Corollary 7.
Let $0<a<1$. Then the operator $a^{-1}C_{\psi_{a}}$ is Li–Yorke chaotic on $H^{2}(\mathbb{C}_{+})$, where
$\psi_{a}(w)=a^{-1}w+(a^{-1}-1).$
Proof.
According to [10], $C_{\varphi_{a}}$ on $H^{2}$ is unitarily equivalent to $a^{-1}C_{\psi_{a}}$ on $H^{2}(\mathbb{C}_{+})$. ∎